Simplity (sqrt7)(sqrt14) my answer 9.89494937
Simpilfy 5(sqrt5)-(sqrt20) my answer 6.708203932 is not right.
Simplify (sqrt5)(sqrt3) my answer 3.87298346
Determine whether (121)^1/3 is rational or irrational. This is the 3rd root of 121.

Found 2 solutions by stanbon, venugopalramana:
Answer by stanbon(75887)

Simplify (sqrt7)(sqrt14)
=sqrt(7^2*2)
=7sqrt2
----------------
Simpilfy 5(sqrt5)-(sqrt20)
=5sqrt5-sqrt(4*5)
=5sqrt5-2sqrt5
=3sqrt5
---------------
Simplify (sqrt5)(sqrt3)
=sqrt15
--------------
Determine whether (121)^1/3 is rational or irrational. This is the 3rd root of 121.
It is irrational because 121 is not a perfect cube.
Cheers,
Stan H.

Answer by venugopalramana(3286)

Simplity (sqrt7)(sqrt14)
=sqrt(7)*sqrt(7*2)=sqrt(7)*sqrt(7)sqrt(2)=7*1.414=9.898
my answer 9.89494937..OK
Simpilfy 5(sqrt5)-(sqrt20)
=5SQRT(5)-SQRT(5*4)=5SQRT(5)- 2SQRT(5)= (5-2)SQRT(5)=3SQRT(5)=6.7082
my answer 6.708203932 is not right.IT IS CORRECT..OK
Simplify (sqrt5)(sqrt3)
=SQRT(15)=3.873
my answer 3.87298346...OK..
Determine whether (121)^1/3 is rational or irrational. This is the 3rd root of 121.
IT IS IRRATIONAL